# A Pre-trained Deep Active Learning Model for Data Deduplication

### v3

@@ 4.2 Analysis of experimental results
For the first part of the experiment, we use all the data in the training set to train the model until convergence, the results obtained are reported in Table 2. PDDM-AL generally outperforms other methods in terms of Precision, Recall and F1 scores. However, in the education dataset, its precision was 0.0483 lower than Dedupe, resulting in a slight lag in F1 score by 0.0065. Nonetheless, PDDM-AL still achieved a higher recall rate by 0.0322 than Dedupe. Compared to Sentencebert, PDDM-AL’s F1 score is consistently higher, with an increase of 0.009, 0.294, 0.246, and 0.056 in respective datasets. Additionally, PDDM-AL also shows improvement in recall rates, which are increased by 0.023, 0.287, 0.267, and 0.192, respectively. Although PDDM-AL’s precision is slightly lower than Sentencebert’s in the Musicbrainz-20-A01 dataset by 0.02, it performs better in the other three datasets with higher precision scores of 0.3013, 0.2237, and 0.1569, respectively.
@@
Model method Musicbrainz-20-A01 GeographicSettlements Education Enterprise personnel
@@
Dedupe 0.9468 0.6594 0.7774 0.9894 0.9257 0.9565 0.9857 0.9088 0.9457 0.782 0.9972 0.8766
Field-based similarity 0.5737 0.6517 0.6102 0.7661 0.8416 0.8021 0.4976 0.7301 0.5918 0.8339 0.86 0.8468
PDDM-AL 0.9583 0.9917 0.9747 0.9687 0.9534 0.961 0.9374 0.941 0.9392 0.994 0.998 0.9979
@@
Figure 7 show the impact of different active learning selection strategies on F1 and Recall, and we compare the random selection strategy with the uncertainty strategy on the three datasets. As can be seen from the figure, with the iterations of active learning, the F1 and Recall of the uncertainty selection strategy are higher than those of the random selection strategy. This is mainly due to the fact that the uncertainty-based selection strategy can select the most uncertain data of the model to be given to the experts for labelling in each round of active learning, and then the model learns from this part of the labelled data to improve its own accuracy.
@@ 5 Conclusion
This study tackles the primary challenges of data deduplication by proposing a pre-trained model based on active learning. It is the first work to introduce active learning at the semantic level to address the problem of deduplication. The model is initially built upon a pre-trained Transformer-based language model, fine-tuned to tackle the deduplication problem as a sequence-to-classification task, which integrates the transformer with active learning into an end-to-end architecture to selectively train the model with the most valuable data. Additionally, we also firstly introduce the R-Drop method with the base model to perform data augmentation on each round of labeled data, reducing manual labeling costs in big data and enhancing the model’s performance with a small amount of labeled data. Experimental results demonstrate that the proposed model outperforms other benchmark methods, highlighting the effectiveness of active learning in data deduplication. Going forward, we aim to explore different selection strategies and combinations of models to further improve the accuracy of the data deduplication model.

### v4

@@ 4.3 Analysis of experimental results
First of all, for the first part of the experiment, we use all the data in the training set to train the model until convergence, the results obtained are reported in Table 2. PDDM-AL generally outperforms other methods in terms of Precision, Recall and F1 scores. However, in the education dataset, its precision was 0.0483 lower than Dedupe, resulting in a slight lag in F1 score by 0.0065. Nonetheless, PDDM-AL still achieved a higher recall rate by 0.0322 than Dedupe. Compared to Sentencebert, PDDM-AL’s F1 score is consistently higher, with an increase of 0.009, 0.294, 0.246, and 0.056 in respective datasets. Additionally, PDDM-AL also shows improvement in recall rates, which are increased by 0.023, 0.287, 0.267, and 0.192, respectively. Although PDDM-AL’s precision is slightly lower than Sentencebert’s in the Musicbrainz-20-A01 dataset by 0.02, it performs better in the other three datasets with higher precision scores of 0.3013, 0.2237, and 0.1569, respectively.
@@
{tblr}
cells = c, cell11 = r=2, cell12 = c=3, cell15 = c=3, cell18 = c=3, cell111 = c=3, hline1,3,7 = -, hline2 = 2-13, Model method & Musicbrainz-20-A01 GeographicSettlements Education Enterprise personnel
@@
Dedupe 0.9468 0.6594 0.7774 0.9894 0.9257 0.9565 0.9857 0.9088 0.9457 0.7820 0.9972 0.8766
Field-based similarity 0.5737 0.6517 0.6102 0.7661 0.8416 0.8021 0.4976 0.7301 0.5918 0.8339 0.8600 0.8468
PDDM-AL 0.9583 0.9917 0.9747 0.9687 0.9534 0.9610 0.9374 0.9410 0.9392 0.9940 0.9980 0.9979
@@
Table 5: F1 for different active learning selection strategies on three datasets
{tblr}
cells = c, cell11 = r=2, cell12 = c=2, cell14 = c=2, cell16 = c=2, hline1,3,8 = -, hline1,3,8 = -, hline2 = 2-7, Number of data pairs & Musicbrainz-20-A01 GeographicSettlements Education
Uncertainty Random selection Uncertainty Random selection Uncertainty Random selection
1000 0.674 0.674 0.798 0.798 0.846 0.846
2000 0.948 0.891 0.936 0.932 0.939 0.890
3000 0.973 0.910 0.961 0.931 0.887 0.834
4000 0.973 0.910 0.961 0.931 0.887 0.834
5000 0.974 0.910 0.961 0.931 0.887 0.834
Table 6: Recall for different active learning selection strategies on three datasets
{tblr}
cells = c, cell11 = r=2, cell12 = c=2, cell14 = c=2, cell16 = c=2, hline1,3,8 = -, hline2 = 2-7, Number of data pairs & Musicbrainz-20-A01 GeographicSettlements Education
Uncertainty Random selection Uncertainty Random selection Uncertainty Random selection
1000 0.970 0.970 0.708 0.708 0.891 0.891
2000 0.924 0.989 0.935 0.929 0.941 0.905
3000 0.989 0.980 0.953 0.942 0.900 0.856
4000 0.989 0.980 0.953 0.942 0.900 0.856
5000 0.992 0.980 0.953 0.942 0.900 0.856
@@
Table 5 and Table 6 show the impact of different active learning selection strategies on F1 and Recall, and we compare the random selection strategy with the uncertainty strategy on the three datasets. As can be seen from the table, with the iterations of active learning, the F1 and Recall of the uncertainty selection strategy are higher than those of the random selection strategy. This is mainly due to the fact that the uncertainty-based selection strategy can select the most uncertain data of the model to be given to the experts for labelling in each round of active learning, and then the model learns from this part of the labelled data to improve its own accuracy. For example, as shown in the figure 8, which illustrates the effect of the two selection strategies on model F1 in the GeographicSettlements dataset. It can be seen that the uncertainty selection strategy can effectively improve the model accuracy, for example, after the third round of active learning, the uncertainty-based selection strategy is 3% higher than the randomly selected selection strategy F1.
@@ 5 Conclusion
This study addresses the key challenges of data deduplication by proposing a pre-trained data deduplication model based on active learning. This is the first work to utilize active learning to address the problem of deduplication at the semantic level. The model is built on a pre-trained Transformer-based language model that has been fine-tuned to solve the deduplication problem as a sequence to classification task. Our proposed model, integrates the transformer with active learning into an end-to-end architecture to select the most valuable data for model training. In addition, we employ the R-Drop method to perform data augmentation on each round of labeled data to reduce the cost of manual labeling in big data and improve the model’s performance with a small amount of labeled data. Experimental results demonstrate that PDDM-AL outperforms other benchmark models, showing the effectiveness of active learning in data deduplication. In the future, we will continue to explore different selection strategies and combinations of different models to further improve the accuracy of the data deduplication model.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Nos. 62276215, 62272398, 62176221, 61976247).